OpenStudy (anonymous):

solve for x: (3x - 1) / 5 - (2x - 4) / 3 = 1 Need help ASAP

OpenStudy (anonymous):

(3x-1)/5-(2x-4)/3=1 cross muliply 3(3x-1)-5(2x-4)=1 9x-3-10x+20=1 -x+17=1 -17 -x=-16 x=16

OpenStudy (anonymous):

OMG Thanks so much!

OpenStudy (anonymous):

no problem lol

OpenStudy (anonymous):

can you tell me how to give out medals? lol